A man is required to pay 57500 pesos at the end of 15 days of 60000 pesos at the end of 60 days. Determine the rate of interest.


1
Expert's answer
2022-01-31T14:22:10-0500

The rate of interest is:

"57500\/(1 + i\u00d715\/365)^{365} = 60000\/(1 + i\u00d760\/365)^{365},"

"(\frac{1 + i×12/73} {1 + i×3/73})^{365} = 60/57.5,"

1.0001166×(1 + 3/73i) = 1 + 12/73i,

i = 0.00095 per annum.